OpenStudy (anonymous):

find the fourier series expansion for the following function f(x)=x sin x for -pi

OpenStudy (unklerhaukus):

can you draw the function?

OpenStudy (unklerhaukus):

can you tell if it is odd even or neither?

OpenStudy (anonymous):

what i know when i test f(-x) is x is odd then sin x is odd when odd n odd it become even right?

OpenStudy (unklerhaukus):

yeah it turns out to be even

OpenStudy (unklerhaukus):

so you only have to find \(a_0\) and , \(a_n\)

OpenStudy (unklerhaukus):

can you set you the \(a_0\) integral?

OpenStudy (anonymous):

what do u mean? is it right i use formula \[2 \div \pi \int\limits_{\pi}^{- \pi} xsinx dx\]

OpenStudy (unklerhaukus):

\[a_0=\frac1l\int\limits_{a}^{a+2l} f(x)\,\text dx\] \[a_0=\frac1\pi\int\limits_{-\pi}^{\pi} x\sin(x)\,\text dx\]

OpenStudy (unklerhaukus):

now we note that the function even, so \[=\frac2\pi\int\limits_0^{\pi} x\sin(x)\,\text dx\]

OpenStudy (unklerhaukus):

(you were close ), can you integrate that?

OpenStudy (anonymous):

i get \[a _{0}=2\] is it correct? then how to find \[a _{n}?\]

OpenStudy (unklerhaukus):

yes! good work

OpenStudy (unklerhaukus):

\[a_n=\frac1l\int\limits_{a}^{a+2l} f(x)\cos\left(\frac{n\pi}lx\right)\,\text dx\]

OpenStudy (unklerhaukus):

its a bit trickier

OpenStudy (unklerhaukus):

(and use the fact its even first then you might need to use some trig formula like \[2\sin(A)\cos(B)=\sin(A+B)+\sin(A-B)\]

OpenStudy (anonymous):

why we used full range equation? even though we know that it is even..